# Models of attractor dynamics in the brain

### v1

@@ Conclusion
Taken together, the four examples reviewed in these lectures highlight the utility of auto-associative neural networks as a powerful and versatile class of analytical models for studying attractor dynamics—a computational motif that recurs across diverse brain regions and cognitive domains. While these models are intentionally simplified and may not capture the full complexity of neural responses to naturalistic stimuli or real-world behavioral tasks, their strength lies in their analytical tractability. By enabling researchers to precisely isolate, manipulate, and interpret specific features of neural dynamics, these models yield mechanistic insights that are often obscured in more complex, high-dimensional systems. Importantly, auto-associative networks provide a conceptually transparent bridge between neural activity and behavior, revealing how stable patterns of activity can support memory, perception, categorization, and decision making. These insights complement the growing class of large-scale, image-computable neuroconnectionist models that prioritize biological realism and scale Doerig et al. (2023).

### v2

@@ Conclusion
Taken together, the four examples reviewed in these lectures highlight the utility of auto-associative neural networks as a powerful and versatile class of analytical models for studying attractor dynamics—a computational motif that recurs across diverse brain regions and cognitive domains. While these models are intentionally simplified and may not capture the full complexity of neural responses to naturalistic stimuli or real-world behavioral tasks, their strength lies in their analytical tractability. By enabling researchers to precisely isolate, manipulate, and interpret specific features of neural dynamics, these models yield mechanistic insights that are often obscured in more complex, high-dimensional systems. Importantly, auto-associative networks provide a conceptually transparent bridge between neural activity and behavior, revealing how stable patterns of activity can support memory, perception, categorization, and decision making. These insights complement the growing class of large-scale, image-computable neuroconnectionist models that prioritize biological realism and scale (Doerig et al., 2023).
@@
Assumptions, limitations, and complementary frameworks. These analytical attractor models make deliberate abstractions (e.g., rate-based units, stylized inputs, simplified noise/adaptation), so they should be read as mechanistic “existence proofs" rather than complete accounts in naturalistic settings. Importantly, “alternative" frameworks often target different levels of description rather than different mechanisms: for example, drift–diffusion models (Bogacz et al., 2006) summarize choice/RT statistics and can be interpreted as a low-dimensional approximation to recurrent decision dynamics (Wang, 2002; Wong and Wang, 2006), while predictive-processing/predictive-coding views emphasize how recurrent feedback can implement inference via prediction-error correction (Barlow, 1993; Friston, 2005). In this sense, these perspectives are typically complementary to—not in competition with—attractor-based circuit mechanisms.